# The Russian-focused embedders’ exploration: ruMTEB benchmark and Russian embedding model design

### v1

@@ A.3.1 Classification
KinopoiskSentimentClassification. In a sentiment classification dataset given a film review, one has to predict whether it is Positive, Neutral, or Negative (3 classes in total). The data was taken from the original dataset Blinov et al. (2013)303030https://huggingface.co/datasets/blinoff/kinopoisk, which contains reviews from July 2004 to November 2012. In the preprocessing phase, we removed all mentions of the final rating from the review texts and balanced the set, leaving only 4,500 samples of each class. The resulting dataset was split into three parts (train, validation, and test), with the class balance preserved.
GeoReviewClassification A classification dataset, where given a review text one has to predict its rating ranging from 1 to 5 (five classes in total). The set is based on the Yandex Maps313131https://yandex.ru/maps reviews323232https://github.com/yandex/geo-reviews-dataset-2023. The original dataset was balanced and split into three parts (train, validation, and test).
@@
RuReviewsClassification A sentiment classification dataset where top-ranked goods from a major e-commerce site were provided, and user-ranked scores were used as class labels on a 5-point scale. The data was sourced from the original dataset RuReviews333333https://github.com/sismetanin/rureviews?tab=readme-ov-file, which contains reviews in the “Women’s Clothes and Accessories” category. During the preprocessing stage, duplicates were removed, and the dataset was balanced, resulting in only 25,000 samples for each class. The resulting dataset was divided into three parts (train, validation, and test) while maintaining class balance.
RuSciBenchGRTNI/OECDClassification. This is a dataset for the classification of scientific text headings. Each article has its OECD and GRNTI headings, with 29 OECD headings and 28 GRNTI headings in the dataset (e.g., Mathematics, Biological Sciences, Economics and Business, etc.). The data was sourced from the original dataset RuSciBench343434https://huggingface.co/datasets/mlsa-iai-msu-lab/ru_sci_bench. During preprocessing, duplicates were removed, the title and abstract were combined, and the set was balanced, leaving only the same number of samples for each class. The resulting dataset was then divided into test and training parts.
InappropriatnessClassification. The dataset aims to predict whether the message is inappropriate or not in the form of binary classification. The data is based on the Inappropriate Messages dataset (version 3)353535https://github.com/s-nlp/inappropriate-sensitive-topics/blob/main/Version3/Inappapropriate_messages.csv Babakov et al. (2021). We binarized the inappropriateness scores using the 0.5 threshold. The resulting dataset was balanced and split into three parts (train, validation, and test), with the class balance preserved.
@@ A.3.3 Multi-Label Classification
SensitiveTopicsClassification. The dataset contains sentences that can be classified into one or more sensitive topics363636https://github.com/s-nlp/inappropriate-sensitive-topics/blob/main/Version3/sensitive_topics.csv Babakov et al. (2021). The original dataset includes 18 classes, all classes are used. Since part of the dataset is not only manually labeled, we first formed a test split from manually labeled examples, and the remaining examples were combined with semi-automatically labeled examples. We have selected the most reliable examples based on the confidence scores indicated in the examples. The final test split consists of 2048 examples and preserves the original class distribution.
@@ A.3.4 Clustering
GeoReviewClustering A clustering dataset based on the Yandex Maps373737https://yandex.ru/maps reviews383838https://github.com/yandex/geo-reviews-dataset-2023, where given a review text one has to cluster the samples according to their rubrics or review categories (e.g., Bank, Supermarket, Pharmacy, etc.). The original dataset was balanced and split into three parts (train, validation, and test). For each review, we took its first rubric as the main label, leaving only samples corresponding to the top 100 most popular labels. This threshold limited the categories exceeding 10,000 examples. The final dataset was converted into the MTEB format.
RuSciBenchGRTNI/OECDClustering. This is a dataset for the clustering of scientific text headings. Each article has its OECD and GRNTI headings, and there are 29 OECD headings and 28 GRNTI headings in the dataset (e.g., Mathematics, Biological Sciences, Economics and Business, etc.). The data was sourced from the original dataset RuSciBench393939https://huggingface.co/datasets/mlsa-iai-msu-lab/ru_sci_bench. During preprocessing, duplicates were removed, the title and abstract were combined, and the set was balanced, leaving only the same number of samples for each class. The resulting dataset was then divided into test and training parts.
@@ A.3.5 Semantic Textual Similarity (STS)
RuSTSBenchmarkSTS. The dataset used for the STS task is derived from the original multilingual STS Benchmark 404040https://github.com/PhilipMay/stsb-multi-mt. This multilingual set comprises various translations of the original English version of the STSbenchmark dataset, with the translations completed using deepl.com 414141https://www.deepl.com/ru/translator. The Russian segment of the dataset was extracted and refined using the RuCoLa Mikhailov et al. (2022) classifier 424242https://huggingface.co/RussianNLP/ruRoBERTa-large-rucola. In all parts of the sets (train/dev/test), instances categorized as not linguistically acceptable were excluded. Additionally, any duplicate entries were eliminated.
@@ A.3.6 Reranking
RuBQReranking. The dataset is based on RuBQ version 2.0 Rybin et al. (2021). The dataset contains examples of questions and paragraphs from Wikipedia. Paragraphs that answer the question are considered relevant. Paragraphs that contain the answer are used as positive documents. Negative documents are paragraphs relevant to the question’s topic but do not contain an answer. We only used questions from the test split with at least nine negative documents. The final test split contains 1551 examples.
@@ A.3.7 Retrieval
RiaNewsRetrieval. The original dataset RussiaSegodnya434343https://github.com/RossiyaSegodnya/ria_news_dataset (also known as RiaNews) consists of news articles and their headlines Gavrilov et al. (2019). Texts are presented in lowercase format, and the capitalization of individual characters has not been changed. Since the article texts are available in HTML, we used the BeautifulSoup444444https://www.crummy.com/software/BeautifulSoup library to clean them of markup. Additionally, texts were normalized, and extra spaces were removed. We also removed, if possible, the first sentence in each article text since it does not relate to the article’s content and is a kind of meta information (“Moscow, 1 Dec — ria news.”). We filtered out the texts of articles with more than 2000 characters so that models limited to a context of 512 tokens could handle the entire text. All examples were deduplicated based on the headline and text of the article. Our final dataset consists of 10,000 randomly sampled headlines as queries, and article texts (724344) are used as documents.
@@ A.5 Detailed Results
Table 8 shows results on individual ruMTEB datasets.
@@
| tiny2 | large-nlu | large-mt-nlu | small | base | large | large-instruct | mistral-7b-instruct | |||
@@
| GeoreviewClassification | 39.6 | 40.0 | 39.7 | 44.7 | 46.1 | 49.7 | 47.2 | 49.7 | 54.2 | 54.4 |
| HeadlineClassification | 74.2 | 79.3 | 77.2 | 73.2 | 75.7 | 75.8 | 70.3 | 78.0 | 86.1 | 86.9 |
| InappropriatenessClassification | 58.6 | 62.5 | 64.6 | 59.2 | 58.8 | 61.6 | 59.9 | 61.3 | 64.8 | 69.5 |
| KinopoiskClassification | 48.2 | 49.5 | 50.3 | 50.0 | 50.9 | 56.6 | 58.2 | 63.3 | 66.0 | 71.7 |
| MassiveIntentClassification | 50.8 | 61.1 | 61.4 | 58.4 | 62.8 | 65.8 | 68.8 | 67.0 | 68.5 | 72.9 |
| MassiveScenarioClassification | 59.1 | 67.6 | 68.1 | 63.9 | 68.2 | 70.9 | 73.4 | 71.8 | 71.6 | 76.3 |
| RuReviewsClassification | 57.0 | 58.3 | 58.3 | 61.2 | 63.0 | 65.3 | 66.9 | 68.0 | 68.1 | 70.1 |
| RuSciBenchGRNTIClassification | 45.7 | 53.9 | 54.2 | 55.0 | 56.3 | 58.2 | 55.8 | 59.3 | 65.0 | 65.7 |
| RuSciBenchOECDClassification | 35.5 | 43.0 | 43.8 | 42.7 | 42.3 | 44.5 | 42.6 | 46.3 | 50.0 | 51.8 |
@@
| GeoreviewClusteringP2P | 41.6 | 59.0 | 58.4 | 58.6 | 54.5 | 60.5 | 63.1 | 65.4 | 74.5 | 76.0 |
| RuSciBenchGRNTIClusteringP2P | 39.1 | 50.4 | 52.2 | 51.1 | 51.6 | 52.0 | 50.2 | 55.5 | 62.2 | 62.0 |
| RuSciBenchOECDClusteringP2P | 35.4 | 46.4 | 47.3 | 44.3 | 44.8 | 45.0 | 43.2 | 47.3 | 53.2 | 54.0 |
@@
| MIRACLRetrieval | 1.9 | 2.0 | 6.2 | 59.0 | 61.6 | 67.3 | 70.1 | 53.9 | 68.0 | 67.8 |
| RiaNewsRetrieval | 13.9 | 11.1 | 21.4 | 70.0 | 70.2 | 80.7 | 83.0 | 78.9 | 83.3 | 78.7 |
| RuBQRetrieval | 10.9 | 12.4 | 29.8 | 68.5 | 69.6 | 74.1 | 71.2 | 66.8 | 74.0 | 75.6 |
@@ A.6 Additional Experimental Findings
Prefixes. We found that E5 prefixes Wang et al. (2022a) performed slightly worse and assume that the variant we use helps to better separate tasks during training. The clustering prefix is more suitable for tasks where thematic identification is required, so in many classification problems, we use it instead of classification, despite the name. We tried adding prefixes with some probability; this improved the results without using prefixes and also worsened the results with them. In addition to stratified sampling, we implemented a sampling strategy that takes pairs with the same prefix but saw no improvement.
@@
Augmentations. Although the model trained on 1500 steps shows comparable results to full training, we tried to apply text level and embedding level augmentations but found no meaningful performance improvement. For the text level, we used character-level augmentation from the Augmentex454545https://github.com/ai-forever/augmentex library Martynov et al. (2024) for both languages. In another experiment, we applied the NEFTune Jain et al. (2023) with 3, 5, and 10 alpha parameters.

### v2

@@ A.3.1 Classification
KinopoiskSentimentClassification. In a sentiment classification dataset given a film review, one has to predict whether it is Positive, Neutral, or Negative (3 classes in total). The data was taken from the original dataset Blinov et al. (2013)333333https://huggingface.co/datasets/blinoff/kinopoisk, which contains reviews from July 2004 to November 2012. In the preprocessing phase, we removed all mentions of the final rating from the review texts and balanced the set, leaving only 4,500 samples of each class. The resulting dataset was split into three parts (train, validation, and test), with the class balance preserved.
GeoReviewClassification. A classification dataset, where given a review text one has to predict its rating ranging from 1 to 5 (five classes in total). The set is based on the Yandex Maps343434https://yandex.ru/maps reviews353535https://github.com/yandex/geo-reviews-dataset-2023. The original dataset was balanced and split into three parts (train, validation, and test).
@@
RuReviewsClassification. A sentiment classification dataset where top-ranked goods from a major e-commerce site were provided, and user-ranked scores were used as class labels on a 5-point scale. The data was sourced from the original dataset RuReviews363636https://github.com/sismetanin/rureviews?tab=readme-ov-file, which contains reviews in the “Women’s Clothes and Accessories” category. During the preprocessing stage, duplicates were removed, and the dataset was balanced, resulting in only 25,000 samples for each class. The resulting dataset was divided into three parts (train, validation, and test) while maintaining class balance.
RuSciBenchGRTNI/OECDClassification. This is a dataset for the classification of scientific text headings. Each article has its OECD and GRNTI headings, with 29 OECD headings and 28 GRNTI headings in the dataset (e.g., Mathematics, Biological Sciences, Economics and Business, etc.). The data was sourced from the original dataset RuSciBench373737https://huggingface.co/datasets/mlsa-iai-msu-lab/ru_sci_bench. During preprocessing, duplicates were removed, the title and abstract were combined, and the set was balanced, leaving only the same number of samples for each class. The resulting dataset was then divided into test and training parts.
InappropriatnessClassification. The dataset aims to predict whether the message is inappropriate or not in the form of binary classification. The data is based on the Inappropriate Messages dataset (version 3)383838https://github.com/s-nlp/inappropriate-sensitive-topics/blob/main/Version3/Inappapropriate_messages.csv Babakov et al. (2021). We binarized the inappropriateness scores using the 0.5 threshold. The resulting dataset was balanced and split into three parts (train, validation, and test), with the class balance preserved.
@@ A.3.3 Multi-Label Classification
SensitiveTopicsClassification. The dataset contains sentences that can be classified into one or more sensitive topics393939https://github.com/s-nlp/inappropriate-sensitive-topics/blob/main/Version3/sensitive_topics.csv Babakov et al. (2021). The original dataset includes 18 classes, all classes are used. Since part of the dataset is not only manually labeled, we first formed a test split from manually labeled examples, and the remaining examples were combined with semi-automatically labeled examples. We have selected the most reliable examples based on the confidence scores indicated in the examples. The final test split consists of 2048 examples and preserves the original class distribution.
@@ A.3.4 Clustering
GeoReviewClustering. A clustering dataset based on the Yandex Maps404040https://yandex.ru/maps reviews414141https://github.com/yandex/geo-reviews-dataset-2023, where given a review text one has to cluster the samples according to their rubrics or review categories (e.g., Bank, Supermarket, Pharmacy, etc.). The original dataset was balanced and split into three parts (train, validation, and test). For each review, we took its first rubric as the main label, leaving only samples corresponding to the top 100 most popular labels. This threshold limited the categories exceeding 10,000 examples. The final dataset was converted into the MTEB format.
RuSciBenchGRTNI/OECDClustering. This is a dataset for the clustering of scientific text headings. Each article has its OECD and GRNTI headings, and there are 29 OECD headings and 28 GRNTI headings in the dataset (e.g., Mathematics, Biological Sciences, Economics and Business, etc.). The data was sourced from the original dataset RuSciBench424242https://huggingface.co/datasets/mlsa-iai-msu-lab/ru_sci_bench. During preprocessing, duplicates were removed, the title and abstract were combined, and the set was balanced, leaving only the same number of samples for each class. The resulting dataset was then divided into test and training parts.
@@ A.3.5 Semantic Textual Similarity (STS)
RuSTSBenchmarkSTS. The dataset used for the STS task is derived from the original multilingual STS Benchmark 434343https://github.com/PhilipMay/stsb-multi-mt. This multilingual set comprises various translations of the original English version of the STSbenchmark dataset, with the translations completed using deepl.com 444444https://www.deepl.com/ru/translator. The Russian segment of the dataset was extracted and refined using the RuCoLa Mikhailov et al. (2022) classifier 454545https://huggingface.co/RussianNLP/ruRoBERTa-large-rucola. In all parts of the sets (train/dev/test), instances categorized as not linguistically acceptable were excluded. Additionally, any duplicate entries were eliminated.
@@ A.3.6 Reranking
RuBQReranking. The dataset is based on RuBQ version 2.0 Rybin et al. (2021). The dataset contains examples of questions and paragraphs from Wikipedia. Paragraphs that answer the question are considered relevant. Paragraphs that contain the answer are used as positive documents. Negative documents are paragraphs relevant to the question’s topic but not the answer. We only used questions from the test split with at least nine negative documents. The final test split contains 1551 examples.
@@ A.3.7 Retrieval
RiaNewsRetrieval. The original dataset RussiaSegodnya464646https://github.com/RossiyaSegodnya/ria_news_dataset (also known as RiaNews) consists of news articles and their headlines Gavrilov et al. (2019). Texts are presented in lowercase format, and the capitalization of individual characters has not been changed. Since the article texts are available in HTML, we used the BeautifulSoup474747https://www.crummy.com/software/BeautifulSoup library to clean them of markup. Additionally, texts were normalized, and extra spaces were removed. We also removed, if possible, the first sentence in each article text since it does not relate to the article’s content and is a kind of meta information (“Moscow, 1 Dec — ria news.”). We filtered out the texts of articles with more than 2000 characters so that models limited to a context of 512 tokens could handle the entire text. All examples were deduplicated based on the headline and text of the article. Our final dataset consists of 10,000 randomly sampled headlines as queries, and article texts (724344) are used as documents.
@@ A.5 Detailed Results
Table 8 shows results on individual ruMTEB datasets. We run evaluation on NVIDIA A100 80GB with torch 2.2.1+cu118 and transformers 4.40.2. Please refer to PR484848https://github.com/embeddings-benchmark/results/pull/19 to access the results.
@@
| tiny2 | large-nlu-ru | large-mt-nlu-ru | small | base | large | large-instruct | mistral-7b-instruct | |||
@@
| GeoreviewClassification | 39.64 | 39.97 | 39.67 | 44.66 | 46.05 | 49.69 | 48.27 | 49.70 | 55.90 | 56.72 |
| HeadlineClassification | 74.19 | 79.26 | 77.19 | 73.94 | 75.64 | 77.19 | 70.32 | 78.00 | 86.18 | 87.02 |
| InappropriatenessClassification | 58.57 | 62.52 | 64.64 | 59.16 | 58.78 | 61.59 | 59.87 | 61.32 | 65.53 | 70.36 |
| KinopoiskClassification | 49.06 | 49.51 | 50.33 | 49.96 | 50.89 | 56.59 | 58.23 | 63.27 | 66.12 | 68.35 |
| MassiveIntentClassification | 50.83 | 61.09 | 61.42 | 58.43 | 62.78 | 65.76 | 68.76 | 66.97 | 67.60 | 73.74 |
| MassiveScenarioClassification | 59.15 | 67.60 | 68.13 | 63.89 | 68.21 | 70.85 | 73.42 | 71.80 | 71.59 | 77.10 |
| RuReviewsClassification | 56.99 | 58.27 | 58.29 | 61.18 | 62.99 | 65.28 | 66.91 | 67.96 | 68.56 | 70.57 |
| RuSciBenchGRNTIClassification | 45.63 | 53.90 | 54.19 | 54.99 | 56.28 | 58.20 | 55.81 | 59.33 | 65.07 | 66.05 |
| RuSciBenchOECDClassification | 35.48 | 43.04 | 43.80 | 41.72 | 42.69 | 43.91 | 42.57 | 46.33 | 50.21 | 52.11 |
@@
| GeoreviewClusteringP2P | 41.58 | 57.12 | 57.07 | 58.57 | 54.46 | 59.59 | 63.09 | 65.42 | 74.34 | 76.32 |
| RuSciBenchGRNTIClusteringP2P | 39.78 | 49.70 | 51.44 | 51.14 | 51.56 | 51.98 | 50.83 | 55.47 | 62.21 | 62.27 |
| RuSciBenchOECDClusteringP2P | 35.98 | 44.48 | 45.36 | 44.33 | 44.79 | 45.12 | 43.21 | 47.29 | 53.09 | 54.13 |
@@
| MIRACLRetrieval | 1.89 | 1.98 | 6.20 | 59.01 | 61.60 | 67.33 | 70.16 | 53.91 | 66.08 | 67.66 |
| RiaNewsRetrieval | 13.92 | 11.11 | 21.40 | 70.00 | 70.24 | 80.67 | 82.99 | 78.86 | 83.26 | 78.94 |
| RuBQRetrieval | 10.87 | 12.45 | 29.80 | 68.53 | 69.58 | 74.13 | 71.22 | 66.77 | 73.90 | 75.98 |
@@ A.6 Additional Experimental Findings
Prefixes. We found that the E5 prefixes Wang et al. (2022a) performed slightly worse and assume that the variant we use helps to better separate tasks during training. The clustering prefix is more suitable for tasks where thematic identification is required, so in many classification problems, we use it instead of classification, despite the name. We tried adding prefixes with some probability; this improved the results without using prefixes and also worsened the results with them. In addition to stratified sampling, we implemented a sampling strategy that takes pairs with the same prefix but saw no improvement.
@@
Augmentations. Although the model trained on 1500 steps shows comparable results to full training, we tried to apply text level and embedding level augmentations but found no meaningful performance improvement. For the text level, we used character-level augmentation from the Augmentex494949https://github.com/ai-forever/augmentex library Martynov et al. (2024) for both languages. In another experiment, we applied the NEFTune Jain et al. (2023) with 3, 5, and 10 alpha parameters.